# The Manifold of the Absolute: Religious Perennialism as Generative Inference Arthur Juliani

### v1

@@ 3. The Generative Ontology
Theological Concept VAE Component Mathematical Role
The Absolute / Godhead Latent Variable $z$ The generative source
Exoteric Dogma / Ritual Observable $x$ The high-dimensional data
Revelation / Tradition Decoder $p_{\theta}(x|z)$ Generates observable forms from $z$
Contemplative Practice Encoder $q_{\phi}(z|x)$ Recovers $z$ from observable forms
Orthodoxy Reconstruction Loss Matching $\phi$ to $\theta$
Self-Transcendence KL Divergence Regularizing posterior to prior
Table 1: Mapping between theological concepts and VAE components. Each element of the religious epistemology finds a structural analog in the variational inference framework.
@@ 3.1 Three Competing Configurations
The Traditionalist school’s central structural claim is that every tradition possesses an exoteric (outer, public, legal-ritual) dimension and an esoteric (inner, initiatic, contemplative) dimension, and that the transcendent unity of religions holds at the esoteric level only [41]. Our framework maps this distinction naturally: the exoteric dimension corresponds to the observable space $\mathcal{X}$ (doctrines, rituals, texts), while the esoteric dimension corresponds to the encoder $q_{\phi}$ and the latent space $\mathcal{Z}$. The unity claimed by perennialism resides in $\mathcal{Z}$, not in $\mathcal{X}$; practitioners converge at the level of contemplative inference, not at the level of observable forms. Not all traditions explicitly recognize this distinction. Protestant Christianity and mainstream Islamic scholarship, for example, reject formal esoteric hierarchies. Yet even these traditions contain contemplative currents (Quaker inner light, Hesychasm, Sufi practice within Sunni orthodoxy) that operationally distinguish between outward observance and inward realization. Our framework requires only this weaker structural claim: that the distinction between observable forms (decoder outputs) and contemplative inference (encoder operation) obtains in practice, even where the tradition’s self-understanding does not frame it in esoteric terms.
[…]
We now evaluate these configurations (Figure 1), alongside syncretism (the combination of elements from different traditions in observable space), through systematic elimination.
@@ 4. The Failure of Exclusivism: Independent Latent Spaces
5. The Failure of Syncretism: The Manifold Hypothesis
Exclusivism fails because it cannot explain convergence. Can one resolve this by combining traditions directly? Syncretism attempts exactly this: the combination or juxtaposition of elements from different traditions in the observable space $\mathcal{X}$. “Christian Yoga,” “Zen Catholicism,” New Age movements that freely mix Native American ritual, Eastern meditation, and Western occultism all combine the outputs of different decoders rather than operating on the latent variable $z$ or on the decoder parameters $\theta$ themselves. The manifold critique developed below shows why this fails geometrically, and it holds under all three configurations (E, U, and P): regardless of whether the latent space is shared or separate, each tradition’s observables lie on a distinct non-linear manifold, and interpolation between manifolds produces incoherent artifacts.
@@ 5. The Failure of Syncretism: The Manifold Hypothesis
5.1 The Geometric Failure of Syncretism
The failure of syncretism is fundamentally a geometric failure (Figure 2). Let $x_{A}$ be a point on the Christian manifold $\mathcal{M}_{A}$ and $x_{B}$ be a point on the Hindu manifold $\mathcal{M}_{B}$. A syncretic approach attempts a linear interpolation:
@@ 5.1 The Geometric Failure of Syncretism
Because the manifolds are non-linear and disjoint (Assumption 5.1 and Claim 5.2), the convex combination $x_{\text{sync}}$ lies off both manifolds for all $\alpha\in(0,1)$.
@@
Figure 2: The failure of syncretism as manifold interpolation. Points $x_{A}$ and $x_{B}$ lie on their respective tradition manifolds, but their linear interpolation $x_{\text{sync}}$ lies in the ambient space, off both manifolds. Such points are “out-of-distribution” and cannot be mapped back to the latent Absolute. (For visual clarity, the figure depicts low-dimensional manifolds in a low-dimensional ambient space. The disjointness argument relies on the high-codimension regime where $\dim(\mathcal{X})$ greatly exceeds $\dim(\mathcal{Z})$.)
To illustrate concretely: consider “Christian Yoga” as an attempted interpolation between Christian contemplative prayer ($x_{A}$) and Hindu asana practice ($x_{B}$). The Christian manifold embeds prayer within a framework of divine personhood, grace, and incarnational theology. The Hindu manifold embeds asana within frameworks of prana, chakras, and non-dual metaphysics. The interpolated practice, which involves performing physical postures while reciting Christian prayers, does not inherit the structural coherence of either tradition. It is, in our terminology, a high-entropy artifact: a point in $\mathcal{X}$ that no valid decoder $p_{\theta}$ would generate.
In generative modeling, this phenomenon is well-known: interpolating between two images in pixel space produces ghostly superimpositions rather than meaningful hybrids [2]. Meaningful interpolation requires operating in the latent space, then decoding through a consistent generative function. The syncretic error is to perform surgery in the observable domain while ignoring the latent structure that gives observations their meaning.
@@ 6.2 Posterior Collapse
What universalism produces is an information-theoretic analog with the same symptom but a different cause. The Data Processing Inequality guarantees that the reduced input $\tilde{x}$ carries less information about $z$ than the full observable $x$, so any encoder trained on $\tilde{x}$ will produce a posterior that is, on average, closer to the prior and less able to disambiguate among competing accounts of $z$. No encoder, however well-optimized, can recover information that the pre-processing step has already destroyed. This makes the universalist failure distinct from standard posterior collapse in an important respect: the latter is an accidental optimization pathology that can be mitigated through better training schedules; the former is a necessary consequence of the input bottleneck. Universalism creates an artificial bottleneck, the “common core,” that is too narrow to carry tradition-specific information, and the resulting inference is systematically degraded: a generic “spirituality” that drifts toward the undifferentiated prior.
It is crucial to distinguish this Shallow Universalism (Configuration U, which seeks unity through feature reduction in $\mathcal{X}$) from Deep Universalism (Configuration P, which locates unity in the shared latent space $\mathcal{Z}$ while preserving tradition-specific decoders). Config U fails because it seeks the universal in the observable space $x$ (where tradition-specific structure is essential) rather than in the latent space $z$ (where, under Config P, unity resides). The distinction is not merely verbal: Config U collapses to a single decoder, discarding the very structure that enables inference, while Config P preserves decoder diversity as the mechanism through which the shared latent source is accessed.
@@ 7.2 Self-Transcendence and the Balance of the ELBO
This interpretation finds support across contemplative traditions. The archetype for this contemplative ascent is Plotinus’s epistrophe (return): the soul progressively strips away multiplicity and sensory particularity until it achieves union with the One [36, VI.9.9–11]. Pseudo-Dionysius inherited this Neoplatonic framework and transmitted it to the Christian mystical tradition [25], where it became the via negativa. The via negativa of Pseudo-Dionysius and Meister Eckhart emphasizes the stripping away of all concepts and images of God until one encounters the “divine darkness,” which is the prior before differentiation. Eckhart’s paradoxical prayer to be rid of the conceptual God expresses KL minimization: even the tradition-specific concept of God must be surrendered to approach the unconditioned Absolute [27]. In Sufi terminology, fanā (annihilation) describes an analogous process: the passing away of the ego-self in the divine presence, followed by baqā (subsistence) in God [38], as the practitioner’s conditioned inference patterns ($q_{\phi}$) are dissolved into the unconditioned reality ($p(z)$).
The Buddhist case is particularly instructive because it reveals a subtlety within the framework. The Madhyamaka teaching of emptiness (śūnyatā) systematically deconstructs the practitioner’s reified concepts, revealing their lack of inherent existence [11]. Strictly speaking, Madhyamaka does not posit a positive ground to “return to”; śūnyatā is a pure negation of inherent existence, mapping naturally to the KL minimization (the stripping away of reified constructs) without entailing a substantive $z$. The notion of “ground luminosity” (prabhasvara) as a positive ultimate ground belongs instead to the Yogācāra and Tathāgatagarbha streams, where ālaya-vijñāna (store-consciousness) and Buddha-nature (tathāgatagarbha) do function as analogs of the prior $p(z)$ [26, 22]. That the same contemplative practices, specifically progressive jhāna absorption and insight meditation, produce convergent phenomenological reports across these doctrinally divergent Buddhist schools is itself evidence for the perennialist reading: the encoder outputs converge even when the traditions’ interpretations of what is encountered differ.
@@ 9. Conclusion
    Syncretism fails geometrically under all configurations. Combining the outputs of distinct decoders in observable space produces off-manifold artifacts. Latent-space interpolation either reduces to orthodox practice within a single tradition (when decoded through one $\theta$) or to tradition genesis (when a new decoder is trained), and in no case yields a valid hybrid inhabiting two existing manifolds (Section 5).
@@
    Perennialism provides the best explanatory fit. Configuration P, a shared latent space with tradition-specific decoders, accounts for both convergence and diversity, and supplies the richest formal account of why committed orthodox practice maximizes the Evidence Lower Bound (Section 7).

### v2

@@ 3. The Generative Ontology
[…]
Decoder $\theta$ Encoder $\phi$
@@
The Traditionalist school’s central structural claim is that every tradition possesses an exoteric (outer, public, legal-ritual) dimension and an esoteric (inner, initiatic, contemplative) dimension, and that the transcendent unity of religions holds at the esoteric level only [41]. Our framework maps this distinction naturally: the exoteric dimension corresponds to the decoder $p_{\theta}$ and the observable space $\mathcal{X}$, while the esoteric dimension corresponds to the encoder $q_{\phi}$ and the latent space $\mathcal{Z}$. Not all traditions explicitly recognize this distinction. Protestant Christianity and mainstream Islamic scholarship, for example, reject formal esoteric hierarchies. Yet even these traditions contain contemplative currents such as the Quaker inner light, Hesychasm, and Sufi practice within Sunni orthodoxy that operationally distinguish between outward observance and inward realization. Our framework requires only this weaker structural claim: that the distinction between observable forms (decoder outputs) and contemplative inference (encoder operation) can be realized in practice.
@@ 3.1 Three Competing Configurations
In addition to these three configurations, we evaluate a fourth approach that operates not on the latent space or decoders but on the observables directly:
Definition 3.8 (Syncretism).
Syncretism combines or interpolates the observable outputs of distinct traditions in $\mathcal{X}$, without modifying the latent space $\mathcal{Z}$ or the decoder parameters $\theta$. Unlike Configurations E, U, and P, syncretism is not a claim about the generative structure of traditions, but an operation on decoder outputs: $x_{\text{sync}}=f(x_{A},x_{B})$ for some mixing function $f$.
We now evaluate these configurations (Figure 2) through systematic elimination.
@@ 4. The Failure of Exclusivism: Independent Latent Spaces
5. The Failure of Syncretism: Non-Linear Decoding
Exclusivism fails because it cannot explain convergence. Can one resolve this by combining traditions directly? Syncretism attempts exactly this: the combination or juxtaposition of elements from different traditions in the observable space $\mathcal{X}$. “Christian Yoga,” “Zen Catholicism,” New Age movements that freely mix Native American ritual, Eastern meditation, and Western occultism all combine the outputs of different decoders rather than operating on the latent variable $z$ or on the decoder parameters $\theta$ themselves. Two critiques, developed below, show why this fails under all three configurations (E, U, and P). First, each tradition’s observables lie on a distinct non-linear manifold in $\mathcal{X}$, and interpolation between manifolds lands off both. Second, and more fundamentally, because the decoders $p_{\theta}$ are non-linear, such interpolated points carry no guarantee of retaining the information about $z$ needed for any encoder to recover the Absolute.
5.1 The Manifold Hypothesis
@@
5.2 The Inferential Failure of Syncretism
The failure of syncretism is ultimately an inferential failure (Figure 3). Let $x_{A}$ be a point on the Christian manifold $\mathcal{M}_{A}$ and $x_{B}$ be a point on the Hindu manifold $\mathcal{M}_{B}$. A syncretic approach attempts a linear interpolation:
@@ 5.2 The Inferential Failure of Syncretism
Because the manifolds are non-linear and disjoint (Assumption 5.1 and Claim 5.2), the convex combination $x_{\text{sync}}$ lies off both manifolds for all $\alpha\in(0,1)$. But the deeper problem is not merely geometric. Each valid observable $x\in\mathcal{M}_{\theta}$ was generated by a specific non-linear decoder $p_{\theta}(x|z)$ and therefore carries structured information about $z$ as refracted through $\theta$: an encoder $q_{\phi}$ can be trained to recover $z$ from such an $x$ because there exists a coherent generative process to invert. For $x_{\text{sync}}$, no such process exists. Because $p_{\theta}$ is non-linear, the blend of two decoders’ outputs does not correspond to decoding through any single $\theta$, and there is no a priori guarantee that the resulting point retains the information about $z$ needed to train an encoder capable of recovering the Absolute. The syncretic practitioner thus faces an unverifiable inference problem: they have constructed an observable with no generative model to invert.
@@
Figure 3: The two failures of syncretism. (a) Manifold interpolation: points $x_{A}$ and $x_{B}$ lie on their respective tradition manifolds, but their linear interpolation $x_{\text{sync}}$ lies off both manifolds in the ambient space $\mathcal{X}$. (b) Non-invertible decoding: each tradition’s observable is generated by a non-linear decoder ($p_{\theta_{A}}$, $p_{\theta_{B}}$) and can be inverted by a matched encoder ($q_{\phi_{A}}$, $q_{\phi_{B}}$) to recover $z^{*}$. The syncretic point $x_{\text{sync}}$ was generated by no decoder, and there is no guarantee that any encoder can recover $z^{*}$ from it.
To illustrate concretely: consider “Christian Yoga” as an attempted interpolation between Christian contemplative prayer ($x_{A}$) and Hindu asana practice ($x_{B}$). Hesychastic prayer was developed to invert the Christian decoder: to recover $z$ from observables embedded in a framework of divine personhood, grace, and incarnational theology. Yogic practice was developed to invert the Hindu decoder: to recover $z$ from observables embedded in prana, chakras, and non-dual metaphysics. The interpolated practice (performing physical postures while reciting Christian prayers) was generated by neither decoder. Because $p_{\theta}$ is non-linear, the practitioner has no way to verify that this hybrid retains the information structure that either tradition’s encoder requires to recover $z$.
In generative modeling, this is well-known: interpolating between images in pixel space produces incoherent superimpositions rather than meaningful hybrids, precisely because the non-linear decoder destroys the correspondence between pixel-space proximity and latent-space proximity [2]. The syncretic error is analogous: performing surgery in the observable domain while ignoring the non-linear generative structure that gives observations their inferential power. A related failure, in which information about $z$ is destroyed not by interpolation but by reduction to common features, is analyzed in Section 6.
@@ 6.2 Posterior Collapse
What universalism produces is an information-theoretic analog with the same symptom but a different cause. The Data Processing Inequality guarantees that the reduced input $\tilde{x}$ carries less information about $z$ than the full observable $x$, so any encoder trained on $\tilde{x}$ will produce a posterior that is, on average, closer to the prior and less able to disambiguate among competing accounts of $z$. No encoder, however well-optimized, can recover information that the pre-processing step has already destroyed. This makes the universalist failure distinct from standard posterior collapse in an important respect: the latter is an accidental optimization pathology that can be mitigated through better training schedules; the former is a necessary consequence of the input bottleneck. Universalism creates an artificial bottleneck, the “common core,” that is too narrow to carry tradition-specific information, and the resulting inference is systematically degraded: a generic “spirituality” that drifts toward the undifferentiated prior. This information-theoretic failure parallels the syncretist failure (Section 5): where syncretism destroys information about $z$ by blending the outputs of incompatible non-linear decoders, universalism destroys it by projecting away the tradition-specific structure through which $z$ is encoded. Both errors operate in $\mathcal{X}$ rather than $\mathcal{Z}$, and both produce observables from which the Absolute cannot be reliably recovered.
@@ 7.2 Self-Transcendence and the Balance of the ELBO
This interpretation finds support across contemplative traditions. The archetype for this contemplative ascent is Plotinus’s epistrophe (return): the soul progressively strips away multiplicity and sensory particularity until it achieves union with the One [36, VI.9.9–11]. Pseudo-Dionysius inherited this Neoplatonic framework and transmitted it to the Christian mystical tradition [25], where it became the via negativa. The via negativa of Pseudo-Dionysius and Meister Eckhart emphasizes the stripping away of all concepts and images of God until one encounters the “divine darkness,” which is the prior before differentiation. Eckhart’s paradoxical prayer to be rid of the conceptual God expresses KL minimization: even the tradition-specific concept of God must be surrendered to approach the unconditioned Absolute [27]. In Sufi terminology, fanā (annihilation) describes an analogous process: the passing away of the ego-self in the divine presence, followed by baqā (subsistence) in God [38], as the practitioner’s conditioned inference patterns ($q_{\phi}$) are dissolved into the unconditioned reality ($p(z)$). The Madhyamaka teaching of emptiness (śūnyatā) systematically deconstructs the practitioner’s reified concepts, revealing their lack of inherent existence [11]. Strictly speaking, Madhyamaka does not posit a positive ground to “return to”; śūnyatā is a pure negation of inherent existence, mapping naturally to the KL minimization (the stripping away of reified constructs) without entailing a substantive $z$. We address this complexity in Section 8.
@@ 9. Conclusion
    Syncretism fails inferentially under all configurations. Combining the outputs of distinct non-linear decoders in observable space produces artifacts that carry no guarantee of retaining the information about $z$ needed for encoder-based recovery of the Absolute. Latent-space interpolation either reduces to orthodox practice within a single tradition (when decoded through one $\theta$) or to tradition genesis (when a new decoder is trained), and in no case yields a valid hybrid inhabiting two existing manifolds (Section 5).
@@
    Perennialism provides the best explanatory fit. Configuration P, a shared latent space with tradition-specific decoders, accounts for both convergence and diversity, and supplies the richest formal account of why committed orthodox practice maximizes the ELBO (Section 7).